# Verifiable cloud-based variational quantum algorithms

### v2

@@ IV.3 Comparisons
Table 1 compares the proposed protocol with related protocols for cloud-based VQAs. Our protocol builds upon the method introduced by Shingu et al. Shingu et al. (2022), incorporating verification to enhance robustness while maintaining low quantum resource requirements. Notably, our protocol requires measurements in specific bases.
@@
An additional advantage of our protocol is its improved tolerance to channel loss. In Shingu et al.’s protocol, the loss of any ancillary qubit during transmission necessitates re-delegating the entire circuit, which is particularly problematic for large circuits. Our protocol, however, allows the server to perform encrypted measurements, achieving tolerance by sharing Bell pairs between the server and client. While both Wang et al.’s Wang et al. (2022) and Li et al.’s Li et al. (2021b) BFK-based protocols also tolerate channel loss, Wang et al.’s protocol requires $6w\cdot d$ qubits with verification, and Li et al.’s protocol requires $16w+$ 1 qubits with verification, where $w$ is the number of qubits in the original NISQ algorithms, and $d$ is the depth of the brickwork state in the BFK protocol. In contrast, our protocol, adapted from Shingu et al.’s approach, requires only $3w$ register qubits and four ancillary qubits. While our protocol employs more qubits due to the added verification step, it still uses significantly fewer qubits than the BFK-based protocols.
V Conclusion and Discussion
@@
In this paper, we propose a protocol for cloud-based VQAs that builds upon the work of Shingu et al., enhancing both verifiability and tolerance to channel loss. We have also demonstrated the blindness and correctness of the protocol, ensuring security and accuracy in cloud-based VQAs. Additionally, we compare the client’s quantum capability requirements and the server’s resource consumption with those in Shingu et al.’s protocol and the BFK-based protocols.
@@
The proposed protocol can be further extended through two potential schemes. First, if the client possesses multiple photon detectors, the $J(\phi)$ operators can be performed on multiple register qubits in parallel, thereby accelerating the protocol’s runtime. However, this approach would increase the ancillary qubit resource consumption. Second, the protocol can be adapted for clients without quantum capabilities by employing the double-server blind quantum computation method Morimae and Fujii (2013b), which is compatible with the proposed protocol.
@@
Further research is necessary in several areas. Reducing communication costs while ensuring security remains a significant challenge. Additionally, exploring alternative verification methods to reduce quantum resource consumption further would be beneficial. Lastly, instead of directly transforming parameterized gate-based quantum circuits into MBQC patterns, adopting an MBQC-native approach Schroeder et al. (2023); Chan et al. (2024); Calderón et al. (2024) for VQAs could offer improved circuit depth reduction, making it suitable for adaptation to cloud-based VQAs. Our protocol has the potential to pave the way for real-world applications of cloud-based quantum machine learning.
Acknowledgements.

### v3

@@ 4.3 Comparisons
Table 1 compares the proposed protocol with related protocols for cloud-based VQAs. Our protocol extends the method introduced by Shingu et al. [30] by incorporating verification and increasing tolerance to channel loss. Regarding the client’s quantum capabilities, our protocol requires only minimal quantum resource consumption; specifically, the client needs only to perform measurements in specific bases, which is significantly easier than preparing qubits in optical systems [9].
@@
An additional advantage of our protocol is its improved tolerance to channel loss. In Shingu et al.’s protocol, the loss of any ancillary qubit during transmission necessitates the re-delegation of the entire circuit, which becomes particularly problematic for large circuits. In contrast, our protocol allows the server to perform encrypted measurements, thereby achieving tolerance by sharing Bell pairs between the server and client. While both Wang et al.’s [28] and Li et al.’s [27] BFK-based protocols also tolerate channel loss, Wang et al.’s protocol requires $6w\cdot d$ qubits with verification, and Li et al.’s protocol requires $16w+$ 1 qubits with verification, where $w$ is the number of qubits in the original NISQ algorithms and $d$ is the depth of the brickwork states in the BFK protocols. In contrast, our protocol, adapted from Shingu et al.’s approach, requires only $3w$ register qubits and four ancillary qubits. Although our protocol employs more qubits due to the added verification step, it still uses significantly fewer qubits than the BFK-based protocols.
5 Conclusion and Discussion
@@
In this paper, we propose a protocol for cloud-based VQAs that extends the work of Shingu et al. by incorporating verifiability and increasing tolerance to channel loss. We have also demonstrated the protocol’s blindness and correctness, ensuring security and accuracy in cloud-based VQAs. Additionally, we compare the client’s quantum capabilities and the server’s resource consumption with those in Shingu et al.’s protocol and BFK-based protocols, demonstrating that our protocol maintains low quantum resource consumption for the server and minimal quantum capabilities for the client.
@@
The proposed protocol can be further extended through two potential schemes. First, if the client possesses multiple photon detectors, the $J(\phi)$ operators can be performed on multiple register qubits in parallel, thereby accelerating the protocol’s runtime. However, this approach increases the server’s quantum resource consumption due to the preparation of additional ancillary qubits. Second, the protocol can be adapted for the client without quantum capabilities by employing the double-server blind quantum computation method [11], which is compatible with the proposed protocol.
@@
Further research is necessary in several areas. Reducing communication costs while ensuring security remains a significant challenge. Additionally, exploring alternative verification methods to reduce quantum resource consumption further would be beneficial. Lastly, instead of directly transforming parameterized gate-based quantum circuits into MBQC patterns, adopting an MBQC-native approach [41, 42, 43] for VQAs could offer improved circuit depth reduction, making it suitable for adaptation to cloud-based VQAs. Our protocol has the potential to pave the way for real-world applications of cloud-based quantum computing.
Acknowledgement
@@
Appendix A Derivation of $J(\phi)$ within Universal Gate Patterns
We analyze the operations within the circuit depicted in Fig. 4(c), involving three ancillary qubits and one register qubit $R=\ket{\psi}_{R}$. Let $A_{1}$, $A_{2}$, and $A_{3}$ represent the ancillary qubits. After the client measures one half of the Bell pair, yielding the result $s_{0}$, the ancillary qubits are prepared in the initial states:
$A_{1}=Z^{s_{0}}R_{Z}\left(\frac{k\pi}{4}\right)\ket{+}_{A_{1}},\quad A_{2}=\ket{+}_{A_{2}},\quad A_{3}=\ket{0}_{A_{3}}.$ (13)
Let $s_{1}$, $s_{2}$, and $s_{3}$ denote the measurement results of $A_{1}$, $A_{2}$, and $A_{3}$, respectively. The resulting operations on the register and ancillary qubits are:
$\begin{split}&\bigl{[}H_{R}\otimes M_{A_{3}}(\phi^{\prime})\bigr{]}CX_{R{A_{3}}}\bigl{[}H_{R}\otimes M_{A_{2}}(Z)\bigr{]}CX_{R{A_{2}}}\bigl{[}H_{R}\otimes M_{A_{1}}(Z)\bigr{]}\\ &\quad CX_{R{A_{1}}}\left[\ket{\psi}_{R}\otimes Z^{s_{0}}R_{Z}\left(\frac{k\pi}{4}\right)\ket{+}_{A_{1}}\otimes\ket{+}_{A_{2}}\otimes\ket{0}_{A_{3}}\right]\\ &=\bigl{[}H_{R}\otimes M_{A_{3}}(\phi^{\prime})\bigr{]}CX_{R{A_{3}}}\bigl{[}H_{R}\otimes M_{A_{2}}(Z)\bigr{]}CX_{R{A_{2}}}\\ &\quad\Bigl{\{}\frac{1}{\sqrt{2}}\left[H_{R}R_{Z}\left(\frac{k\pi}{4}+s_{0}\pi\right)\ket{\psi}_{R}\otimes\ket{0}_{A_{1}}\right]\otimes\ket{+}_{A_{2}}\otimes\ket{0}_{A_{3}}+\\ &\quad\frac{1}{\sqrt{2}}\left[H_{R}R_{Z}\left(-\frac{k\pi}{4}-s_{0}\pi\right)\ket{\psi}_{R}\otimes\ket{1}_{A_{1}}\right]\otimes\ket{+}_{A_{2}}\otimes\ket{0}_{A_{3}}\Bigr{\}}.\end{split}$ (14)
The measurement result $s_{1}$ of $A_{1}$ determines the $Z$-rotation angle for |$\ket{\psi}_{R}$. We can simplify the above expression by omitting $A_{1}$:
$\begin{split}&\left[H_{R}\otimes M_{A_{3}}(\phi^{\prime})\right]CX_{R{A_{3}}}\left[H_{R}\otimes M_{A_{2}}(Z)\right]CX_{R{A_{2}}}\\ &\quad\left[H_{R}R_{Z}\left((-1)^{s_{1}}\left(\frac{k\pi}{4}+s_{0}\pi\right)\right)\ket{\psi}_{R}\otimes\ket{+}_{A_{2}}\otimes\ket{0}_{A_{3}}\right]\\ &=\left[H_{R}\otimes M_{A_{3}}(\phi^{\prime})\right]CX_{R{A_{3}}}\left[I_{R}\otimes\left(M_{A_{2}}(Z)H_{A_{2}}\right)\right]CZ_{R{A_{2}}}\\ &\quad\left[R_{Z}\left((-1)^{s_{1}}\left(\frac{k\pi}{4}+s_{0}\pi\right)\right)\ket{\psi}_{R}\otimes\ket{0}_{A_{2}}\otimes\ket{0}_{A_{3}}\right].\end{split}$ (15)
Since the $CZ$ operation does not entangle |$\ket{0}_{A_{2}}$ with the register qubit, we can omit the operations on $A_{2}$. Thus, we obtain the following expression for the remaining qubits:
$\begin{split}&\left[H_{R}\otimes M_{A_{3}}(\phi^{\prime})\right]CX_{R{A_{3}}}\\ &\quad\left[R_{Z}\left((-1)^{s_{1}}\left(\frac{k\pi}{4}+s_{0}\pi\right)\right)\ket{\psi}_{R}\otimes\ket{0}_{A_{3}}\right]\\ &=\frac{1}{\sqrt{2}}H_{R}R_{Z}\left((-1)^{s_{1}}\left(\frac{k\pi}{4}+s_{0}\pi\right)-\phi^{\prime}\right)\ket{\psi}_{R}\otimes\ket{0}_{A_{3}}+\\ &\quad\frac{1}{\sqrt{2}}XH_{R}R_{Z}\left((-1)^{s_{1}}\left(\frac{k\pi}{4}+s_{0}\pi\right)-\phi^{\prime}\right)\ket{\psi}_{R}\otimes\ket{1}_{A_{3}}.\end{split}$ (16)
The measurement result $s_{3}$ of $A_{3}$ determines the state of $R$:
$R=X^{s_{3}}H_{R}R_{Z}\left((-1)^{s_{1}}\left(\frac{k\pi}{4}+s_{0}\pi\right)-\phi^{\prime}\right)\ket{\psi}_{R}.$ (17)
Applying Eq. 6, the above simplifies to:
$\begin{split}R&=X^{s_{3}}H_{R}R_{Z}\left(\phi-r\pi\right)\ket{\psi}_{R}\\ &=X^{s_{3}\oplus r}H_{R}R_{Z}(\phi)\ket{\psi}_{R}.\end{split}$ (18)
Thus, $J(\phi)$ is obtained with Pauli byproducts on the register qubit $R$.